HUBBLE SPACE TELESCOPE DAILY REPORT #1159 PERIOD COVERED: 0700 EST Daily Status Report as of 165/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/BL 5656 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats Part 2) The Faint Object Spectrograph (Blue side) was used to make calibration observations of star HD207757 (aka, AG Pegasi) and object NGC-188-225. A total of seven observations was received, and have not yet been completely analyzed. However, some of the target acquisition data was received in real-time. The peakups had maximum counts in groups 1 and 3, but the counts were much lower than the previous set using the old flight software and appeared to have missed the target. At the present time no explanation has been found. 1.2 Completed GHRS/1 5594 (GHRS Side 1 Sensitivity Monitor: Cycle 4) The High Resolution Spectrograph (Side One) was used to make calibration observations of star BD28D4211 to monitor the long-term sensitivity of the system. A total of fourteen observations was received, but not displayed due to lack of time. The target acquisition was seen in real time and the target was in map 5 of the 5X5 spiral search, with a target search flux of 7135 counts. The Small Science Aperture peakup (5X5) found the target at the eighth dwell point with a target flux of 4259 counts. No anomalies were noted. 1.3 Completed WF/PC-2 5568 (WF/PC-2 Cycle 4 Calibration: Decon) The WF/PC-2 was used to make internal flat calibration observations, K-Spots, and darks as part of a decontamination operation. Six observations were received, but there was not enough time to display them. No anomalies were noted. 1.4 Completed FOC/96 5526 (FOC UV Throughput Monitoring: Cycle 4) The Faint Object Camera (f/96) was used to make calibration observations of BPM16274. The first observations were received and displayed in real-time. A corrective pointing maneuver was successfully calculated, and uplinked to the spacecraft. The target was placed at the fiducial point of the aperture. No anomalies were noted. 1.5 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 5563 (WF/PC-2 Cycle 4 Photometric Calibration Monitor 1: UV/OPT STD) The WF/PC-2 was used to make photometric calibration images of ultraviolet standard star GRW+70D582 and monitor the photometric stability and quantum efficiency of the WF/PC-2. A Loss of Lock (LOL) occurred during the observations (See Operations Note 2.4). The observational data has not yet been received. The DOC is working on a Line Outage Recorder (LOR) playback. Analysis is pending the receipt of the data. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 10 Successful : 10 A Loss of Lock (LOL) occurred at 165/0938Z near a Day/Night transition. The Take Data Flag (TDF) dropped and science data observations were lost. Subsequence lock operations were successful. 2.2 FHST Updates: Scheduled :11 Successful :11 2.3 Significant HSTARs: None 2.4 Operations Notes: The HST observatory experienced a quiet day as each subsystem reported nominal performance. 3.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-